OpenStudy (anonymous):

How do I solve cos(2arcsin 2x) with a triangle? (Or just in general how to solve it.

OpenStudy (ash2326):

\[\cos (2\sin^{-1} 2x)\] DO you know the range of \(\sin^{-1} x\)?

OpenStudy (phi):

Here's how I would do it. t= arcsin(2x) means there is an angle t such that sin(t)= 2x  so your problem is cos(2t)= cos^2(t)- sin^2(t) use the drawing to figure out cos(t)